### v1

# Steering AI-Driven Personalization of Scientific Text for General Audiences

Abstract.
Digital media platforms (e.g., social media, science blogs) offer opportunities to communicate scientific content to general audiences at scale. However, these audiences vary in their scientific expertise, literacy levels, and personal backgrounds, making effective science communication challenging. To address this challenge, we designed TranSlider, an AI-powered tool that generates personalized translations of scientific text based on individual user profiles (e.g., hobbies, location, and education). Our tool features an interactive slider that allows users to steer the degree of personalization from 0 (weakly relatable) to 100 (strongly relatable), leveraging LLMs to generate the translations with given degrees. Through an exploratory study with 15 participants, we investigated both the utility of these AI-personalized translations and how interactive reading features influenced users’ understanding and reading experiences. We found that participants who preferred higher degrees of personalization appreciated the relatable and contextual translations, while those who preferred lower degrees valued concise translations with subtle contextualization. Furthermore, participants reported the compounding effect of multiple translations on their understanding of scientific content. Given these findings, we discuss several implications of AI-personalized translation tools in facilitating communication in collaborative contexts.
AI Personalization, Generative AI, Large language models, Science communication, Steering AI
ccs: Human-centered computing Interactive systems and toolsccs: Computing methodologies Natural language generation
1. Introduction
Science communication refers to the process of “communicating complex scientific information with general audiences to improve public awareness, interest, and understanding of science” (Burns et al., 2003). Both academic and industry professionals strive to make their ideas more accessible. As a result, science communication platforms, such as science blogs (Hogan, 2024) and science magazines (Times, 2024), have been translating scientific information into more accessible formats for decades. However, a single version of a text often falls short of meeting the diverse needs of a general audience (August et al., 2024). For instance, a journalist with an interest in quantum physics would have vastly different context and comprehension needs than a college student majoring in physics.
Scientific research articles often employ domain-specific language, including jargon and complex sentence structures, which can pose significant barriers to comprehension for general audiences. Large language models (LLMs) offer the capability to personalize content to various user contexts, transforming both the style (e.g.,  casual to formal (Das et al., 2023)) and content (e.g.,  simplifying complex ideas (August et al., 2024)). This adaptive personalization creates opportunities for enhancing audience engagement and understanding through more effective content dissemination (Kim et al., 2024). For example, LLMs can tailor the tone and style of complex concepts like scientific information to align with individual language preferences, thereby improving readability and comprehension across various fields (August et al., 2023; Das et al., 2023; Ding et al., 2023). Notably, AI can achieve this personalization at scale (i.e.,  AI-scalable personalization), generating multiple tailored variations of content for diverse audiences (August et al., 2024; Riedl, 2010; Kim et al., 2024). This scalability positions LLMs as powerful tools in advancing inclusive and effective science communication (Razack et al., 2021).
Among the many strategies in science communication, analogies have proven effective in translating technical content into more accessible forms (August et al., 2020). For instance, the structure of the solar system is often used as an analogy to explain the structure of an atom. While such analogies can broaden understanding of scientific information, they are typically generalized and may not resonate equally well with all audiences, as individual comprehension is shaped by societal, cultural, educational, and personal backgrounds. For example, the space analogy assumes familiarity with the solar system (Kim et al., 2024). LLMs have demonstrated the ability to generate personalized analogies that can help people understand complex concepts and ideas within their own context (Ding et al., 2023). Building on this capability, we leverage LLMs to generate multiple personalized analogies tailored to individual readers’ contexts and comprehension levels.
We designed and implemented TranSlider (Translate through Slider), an interactive reading interface that enables users to steer the degree of personalization in scientific text through an adjustable slider. TranSlider allows the user to specify a degree of personalization from 0 to 100 and utilizes their background information (e.g., education, hobbies, location) to present relevant analogies. The slider enables intuitive exploration of various personalization degrees, allowing users to quickly review multiple translations. We employed TranSlider as a research probe to ask the following research questions:
  • RQ1:
    What is the utility of AI-driven personalized translations of scientific text?
  • RQ2:
    What is the impact of interactive reading features on user experience? Specifically:
    • 2.1:
      How does exploring multiple translations influence readers’ comprehension and engagement with scientific text?
    • 2.2:
      How does the slider interaction to steer the degree of personalization influence readers’ comprehension and engagement with scientific text?
To answer these questions, we conducted a user study with 15 non-expert participants who used the tool to understand two scientific texts by exploring multiple personalized translations. We conducted post-session semi-structured interviews to elicit feedback.
Participants generally found the analogies useful in understanding the content. Some favored general analogies (e.g., a participant liked a construction analogy describing the body’s cells as building sites—low personalization), while others preferred detailed, personalized analogies (e.g., a participant appreciated a baking analogy to explain harmful emissions from lithium-ion batteries—high personalization). Participants noted that reading multiple translations with varied analogies allowed them to piece together a fuller understanding of the scientific text, correcting misunderstandings along the way. Overall, users found the tool beneficial for learning about unfamiliar topics but were cautious about the reliability of AI-generated content. Based on these findings, we discuss the need for HCI techniques to steer models towards human pluralistic preferences, rather than solely relying on machine learning methods that assume uniformity in human experiences.
In summary, our study has the following three contributions:
  1. (1)
    A novel slider-based interaction to enhance understanding of scientific content through the exploration of analogy-driven personalized translations
  2. (2)
    An investigation to examine the utility, benefits, and limitations of personalized translations in science communication
  3. (3)
    Implications for designing such AI-personalized translation tools and broader applications in collaborative and education contexts such as facilitating cross-disciplinary communication
2. Related Work
Below, we review prior work covering science communication on digital media platforms and personalization of language content for better science communication.
2.1. Science Communication on Digital Media Platforms
Science communication takes many forms in modern media for communicating complex scientific information to general audiences for improving public awareness, interest, and understanding of science (Burns et al., 2003). Traditional channels include museums (Bell, 2008), exhibitions (Long et al., 2021), and TV series  (Sagan et al., 1980), while web-based digital platforms have emerged as a powerful new stream for the broader public. For example, online videos have become an influential source for learning science, especially during the COVID-19 pandemic, when many people turned to digital resources for scientific information (Breslyn and Green, 2022). Social media platforms such as Reddit (e.g., r/science), Twitter (now X), Bluesky, and Mastodon have evolved into vibrant spaces where some scientists actively share their research, promote their publications, and engage in public discussions (Gero et al., 2021; Jones et al., 2019). Additionally, science blogs have been valuable channels for science communication, offering tailored narratives of complex scientific content for general audiences (Jarreau and Porter, 2018). One of the core benefits of science communication on these digital media is their scalability in reaching broader audiences (e.g.,  (Nguyen et al., 2024; Jones et al., 2019)). Recent studies show that more than half of college students actively use blogs as a learning resource (Head et al., 2017).
Prior HCI/CSCW literature has uncovered several key challenges that scientists face in science communication: engagement, translation, and dissemination (Williams et al., 2022). First, scientists rarely participate in digital media platforms (Zhu and Purdam, 2017). This limited engagement may be associated with multiple factors, including conventional academic culture that has historically put scientists as neutral observers rather than active communicators (Martinez-Conde, 2016; Peters, 2013), and practical constraints such as time limitations and lack of communication training (Besley et al., 2015). While recent research shows a gradual shift towards more active public engagement by scientists (Brüggemann et al., 2020), the extent and impact of this change remain to be fully understood. Secondly, even for those scientists who engage with digital communication, translating complex scientific concepts into accessible content for general audiences is challenging (Rice and Giles, 2017). These challenges may be compounded by the varied characteristics of general audiences in digital media platforms, where readers have widely differing backgrounds, levels of scientific literacy, and preferences (Schäfer, 2017; Su et al., 2014; Rice and Giles, 2017). A single version of scientific content translation would not sufficiently serve this wide spectrum of general audiences, yet creating multiple versions tailored to different expertise levels would also be infeasible for individual scientists (Bucchi and Trench, 2014; Fischhoff, 2013). Lastly, such translation challenges can lead to broader concerns about dissemination, particularly regarding the potential for miscommunication and misinformation on social media platforms (Cook et al., 2018). Scientists may be hesitant to engage in digital communication because the dissemination of inaccurate interpretations or misrepresentations of their work could have unwanted consequences for public understanding and their professional reputation (Wu et al., 2019; Loroño-Leturiondo and Davies, 2018).
Given these challenges, our study focuses on addressing the translation barrier in personalized science communication on digital media platforms. Specifically, we explore how to facilitate creating multiple translations of scientific content to effectively accommodate large-scale general audiences with varying levels of scientific background.
2.2. Personalization of Language Content for Better Science Communication
Prior research has explored various approaches to make complex science texts more accessible and comprehensible to broader audiences, from using plain language summaries (August et al., 2023) and language constructs like analogies (Kim et al., 2022) to using interactive techniques and systems (Chang et al., 2023; Fok et al., 2023). Across these works, limited explorations have been made in personalization of scientific translations to facilitate better understanding, which this paper explores.
Personalization has been explored across various forms of language content, including news articles (Adar et al., 2017), social media posts (Zhang et al., 2022), and novels (Kim et al., 2024), to better accommodate individual users’ needs and preferences. Prior research has demonstrated the effectiveness of such personalization across several domains (e.g., healthcare, journalism) through approaches such as using patients’ demographic and medical information to tailor content  (Dimarco et al., 2006), using users’ location data to make news content relevant (Adar et al., 2017), and using analogies to make measurements (e.g., distances, areas) comprehensible to a user’s context (Kim et al., 2016). For instance, employing familiar objects as an analogy (e.g., 28 lbs is about the weight of a microwave) to personalize the unit expressions (Hullman et al., 2018). In science communication contexts, past work has shown AI’s potential in large scale translation and dissemination through generating personalized summaries of scientific texts for general audiences (August et al., 2024). In this paper, we build on these ideas to explore using LLMs for generating and customizing scientific text explanations using user provided data for analogies.
While relying on users’ contextual information (e.g., location) delivers personalized outcomes, these systems typically operate autonomously making personalization decisions without the users’ direct input or control. This raises several potential concerns. First, there are privacy issues, as these systems require users to share personal data, often assuming their willingness to do so or even collecting such information without notice (implicitly) (e.g.,  location information based on IP address (Ma et al., 2024), personal profile based on search engine account (Hannak et al., 2013; Majumder and Shrivastava, 2013)). Second, users generally lack control over which specific data they share for personalization purposes. Additionally, individual users may have different preferences regarding how much of their profile information should be integrated with the tailored content. To address these issues in user agency and privacy control, we designed a steerable reading interface that enables users to control the degree of personalization with adjustable sliders as well to edit profile information directly for personalizing the translations to their context.
3. TranSlider: Design and Implementation
We designed and implemented TranSlider (Translate through Slider), an AI-powered tool to help individuals with no scientific backgrounds better understand scientific texts by personalizing them to their context using a slider interaction and user profiles.
Figure 1. The interface design of TranSlider. (A) The AI-personalized title of the science blog post appears here. (B) AI-personalized translations of science articles will appear here. (C) Users can navigate the slider to adjust their preferred degree for AI personalization of science blog intro. (D) For each personalized translation, a history item is generated and stacked in the history box. Users can revisit any previous history items to see the translations. (E) Users can click the edit profile button to see the Profile Edit window [II] to adjust their profile information. (F) The Finish button triggers a pop-up [III] to collect the user’s interest in reading the original paper. We use the response as a proxy measure for their interest in the content after reading the personalized translations.
3.1. Interface Design
The interface comprised three main components: the translation panel, the history box, and the edit profile dialog box (Figure 1). The header provided instructions on using the interface; these instructions were brief as participants received a detailed tutorial at the start of the study.
3.1.1. Translation Panel
The translation panel housed the main personalization features (Figure 1 A, B, C). At the top, the personalized title of the scientific article being viewed is displayed, with a box below presenting the personalized translation. A slider, labeled from weak to strong personalization, allowed participants to adjust the degree of personalization. The current personalization degree, a numeric value between 0 and 100, was shown just beneath the slider. Participants were instructed to adjust the slider to their preferred degree value and click Translate at the bottom. This action triggered a call to an LLM, which generated and displayed a personalized translation in the box above.
3.1.2. History Box
The history box on the right allowed participants to revisit, review, and compare translations they had seen during the session. Each new translation was automatically added to the stack, labeled with its personalization degree and the time of generation. Participants could click on any past translation to view it, which would update the translation panel accordingly. They could also mark a translation as a favorite by clicking on the star icon beside it. Additionally, participants could delete translations from the history by clicking the cross button, though we found that participants rarely used this feature in our study.
3.1.3. Edit Profile Dialog Box
The edit profile dialog box was accessible by clicking a button on the page header (Figure 1E). This displayed a short form (Figure 1[II]) containing information about the participant including their background, age, hobbies, location, and favorite food. This profile information was used to tailor the translation to the participant’s context. The study consisted of two sessions: in the first, participants could not see or edit their profiles, while in the second, they were allowed to update their profiles to further tailor the personalization.
3.1.4. Finishing a Session
Once participants were satisfied with their understanding of the scientific material, they were instructed to click the finish button (Figure 1F) in the top right corner. This action triggered a pop-up (Figure 1[III]) that displayed a thank you message and an additional question asking if they would like to receive the original research paper PDF via email after the study. This served as a proxy measure for their interest in the content after reading the personalized translations.
3.2. Implementation and Prompt Design
We built the interface using React, a popular JavaScript library for front-end development, and implemented the backend in JavaScript. When a participant requested a translation at a specific personalization degree, an API call was made to OpenAI’s GPT-4o model to generate the translation. As shown in Figure 2, we used a detailed prompt template to guide the model’s output.
Figure 2. Prompt design for generating personalized translations. The prompts (A, B, C, and D) are combined in order, and then sent to LLM through an API call to generate tailored translations of scientific articles.
3.2.1. LLM Prompt
The prompt began with a high-level task overview, instructing the model to use analogies to explain a science article to general audiences. It then defined the personalization spectrum, providing clear descriptions for degrees 0, 50, and 100. Following this, the prompt included in-context examples of both generic and personalized translations. To improve quality, we applied chain-of-thought prompting, guiding the model to reason through its responses. Finally, the model was given the scientific article and asked to produce a two-paragraph translation while avoiding titles, honorifics, or other formalities. The exact prompt template is given in Figure 2.
We arrived at this prompt template through iterative experimentation. Once the research team was satisfied with the LLM’s output, we conducted a pilot study with three research scientists who had published their work (in Computer Science, Biomedical Engineering, and Electrical Engineering, respectively) in the past three years to gather feedback on personalization quality and to confirm that the model was not hallucinating content in its translations. We used their own research papers in this pilot study to let them verify the personalized translations. We proceeded with the main study after receiving this confirmation from all three of them.
4. User Study
Our study aimed to obtain readers’ feedback and usage data regarding two aspects of the tool: i) the utility of AI-personalized translation of scientific text (RQ1), and ii) the impact of interactive reading features on user experience (RQ2). The study consisted of two sessions, designed to gradually introduce tool features to participants. In the first session, participants interacted with TranSlider that implicitly personalizes scientific articles based on their user profiles, similar to how web browsers and social media platforms display personalized content based on collected user metadata (e.g., Google Chrome’s personalization through Google accounts). The second session introduced a more full-featured interface where participants have the full autonomy to explicitly edit their profiles to manipulate the direction of personalization. Throughout both sessions, participants control a slider to explore different degrees of personalization along a continuous spectrum. We describe the detailed study procedure in Section 4.3.
4.1. Science Articles
We selected scientific articles from two distinct domains: health and environment. While health-related topics can generate strong interest among broad audiences due to their direct personal relevance (e.g.,  (Jacobs et al., 2014; De Choudhury et al., 2014; Schorch et al., 2016)), environmental topics might feel less immediately relevant to individuals’ personal contexts. We deliberately chose these contrasting domains to demonstrate that the utility of AI-personalized translations may not be dependent on a particular topic; if participants could perceive the value of AI-personalized translations across both more personal (health) and less personal (environmental) topics.
To select science articles that the general audience would be interested in, we searched articles in the Reddit sub-channel r/science [^1] https://www.reddit.com/r/science/. The first author compiled a list of 10 candidate articles for each domain from top-rated posts within the past year. Through iterative discussions among collaborators, we narrowed our selection to two articles: one in health (Kong et al., 2024) and one in environment (Guelfo et al., 2024). To mitigate potential ordering effects, we counterbalanced the presentation sequence across participants, with half viewing the health article first followed by the environment article, and vice versa for the remaining participants (last two columns in Table 1).
4.2. Participants
Experience Science Articles
Gender Age Education w/ Sciα w/ LLMsβ Session I Session II
A P1 Female 43 UX Design 3 Yes — No health environment
P2 Male 23 Economics 2 Yes — No health environment
P3 Male 32 Industrial Eng. 4 Yes — Yes health environment
P4 Female 26 Industrial Eng. 2 Yes — No health environment
P5 Female 24 Statistics 2 Yes — No health environment
P6 Male 34 Business 1 Yes — Yes health environment
P7 Female 24 Business 4 Yes — No health environment
[…]
P9 Female 25 Economics 2 Yes — No environment health
P10 Female 23 Biochemistry 4 Yes — Yes environment health
P11 Female 24 Computer Sci. 2 Yes — No environment health
P12 Female 23 Information Sys. 3 Yes — No environment health
P13 Female 25 Business 5 Yes — No environment health
P14 Male 24 Cognitive Sci. 3 Yes — Yes environment health
P15 Female 25 Business 2 Yes — No environment health
Table 1. Demographic and background of participants. P1 does not have a degree, but has a certificate in UX Design. αExperience reading scientific text in the last three years (1: Never, 2: A few times per year, 3: A few times per month, 4: A few times per week, 5: Daily). βExperience with LLMs (first value represents general LLMs experience, second value represents LLM use specifically for reading).
We recruited 15 participants (10 female and five male) from our organization through internal mailing lists and word-of-mouth referrals. While our organization employs highly educated individuals with an analytical mindset, we used a diverse recruitment strategy to broaden the applicability of our study. Specifically, to ensure that our sample represented general readers rather than scientists or science writers, we excluded individuals with a doctoral degree or equivalent job titles (e.g., research scientists). All participants were fluent in English and aged 18 or above (range: 23-43, M=26.7, SD=5.3).
Participants came from diverse educational backgrounds, including Business, Economics, Industrial Engineering, Computer Science, Information Systems, Cognitive Science, Biology, and Statistics. We asked about their frequency of reading scientific articles (e.g., science blogs) and experience with LLMs (e.g., ChatGPT). Nine participants reported actively reading scientific texts at least once a month, while three had not engaged with any scientific articles in the past three years. All participants had prior experience with LLMs, with eight actively using them to comprehend complex texts like technical reports or academic papers. However, four participants were familiar with LLMs but not actively utilizing them daily. To further characterize our sample, we collected information on participants’ highest level of education. Table 1 provides detailed demographic data.
4.3. Procedure
The study was conducted 1-on-1 and in-person. We audio and screen recorded study sessions that took an hour on average (M: 60.5 min, SD: 10.9 min). All participants received a 75 USD virtual gift card as compensation for their time.
4.3.1. Pre-study Survey (5 min)
Before conducting the exploratory study, we collected participant background information through a pre-study survey. This included their experience with LLMs and reading scientific texts. We also gathered personal information such as their professional background, hobbies, familiar locations, and favorite foods, which was later used for implicit AI personalization in Session I.
4.3.2. Tutorial (15 min)
After the pre-study survey, we showed the tool interface through a shared monitor in the same room. We explained the tool’s purpose and provided a tutorial using a structured script. The tutorial covered tool features such as sliders, buttons, history box items, and included a demonstration of an example interaction. Throughout the tutorial, we encouraged participants to ask questions to ensure their complete understanding of the tool. After the guided tutorial, participants freely explored the tool until they felt comfortable to begin the main study sessions.
4.3.3. Session I (Implicit) & II (Explicit) (10 min $\times$ 2)
As described earlier, all participants conducted the two study sessions in the same order to ensure gradual exposure to the full-featured tool interface. In the first session, participants interacted with the tool version that implicitly employs their profile to tailor the scientific articles. We explained to participants that the tool used the profile data they provided in the pre-study survey for personalization. To illustrate this concept, we gave them familiar examples of personalization, such as how Google Chrome leverages users’ account data or how social media platforms provide personalized content based on their accounts (Jhaver et al., 2023). After completing the first session, they began the second session with our fully featured tool that enabled users to explicitly edit their profiles.
In both sessions, we asked participants to freely explore varying degrees of personalization through the slider control. We ensured they reviewed at least five different degrees of personalized translations, and then selected three favorite translations by marking a star to the associated history item. After completing each session, participants responded to a post-session survey, where they rated their familiarity with the scientific articles and wrote the key takeaways about the translated article in 1-2 sentences. We then conducted a post-session interview focusing on: (a) their thought process while interacting with the personalization slider, (b) their rationale for selecting favorite translations, and (c) examples of translations they disliked and their reasons.
4.3.4. Debrief Interview (20 min)
After completing both sessions (implicit and explicit), we conducted a semi-structured debrief interview aligned with our research questions. The questions included i) the participants’ perceived quality and utility of AI-personalized translations (RQ1), ii) their experience with interactive reading features such as multiple translations, history items, and the slider (RQ2), iii) any particular aspects they liked or disliked, and iv) overall concerns and suggestions for improvement. At the end, participants were free to ask any questions or share additional comments.
4.4. Analysis
4.4.1. User Behavioral Logs
We recorded user logs to analyze participants’ tool exploration behaviors. The collected data includes user profile information (implicit), profile edits (explicit), explored degrees of personalization, all generated translation texts, favorite translations marked by users, deleted translations, and precise timestamps for all these actions. Then we analyzed them by plotting and computing descriptive statistics (e.g., mean, std, max, and min). We did not conduct a statistical analysis of our data between conditions or groups, because we had a small sample size ($N=$15) (Gelman and Carlin, 2014). More importantly, our study design was intended to enable qualitative analysis.
4.4.2. Interview Logs
All interview recordings were transcribed using automated transcription services. We conducted thematic analysis (Braun and Clarke, 2006) over the transcribed quotes that were organized by interview questions for each participant. For the open-coding process, the first author read quotes and labeled them with initial codes (i.e., low-level codes) using Miro board cards [^2] https://miro.com/. Then the first and the second authors collaboratively reviewed the low-level codes to discuss and group codes into potential themes (i.e., high-level themes). The first author structured a thematic map showing themes for each corresponding research question. Next, the research team collaborated to refine the thematic map more clearly, build a coherent narrative, and reach a consensus. During this process, we excluded themes that were not directly relevant to our research questions. Ultimately, we finalized the themes for each research question: the utility of AI-personalized translations (RQ1), the impact of interactive reading features on user experience (RQ2), and proposed use cases and associated concerns.
5. Findings
We first set the context by summarizing user behavior using log data from TranSlider. Then we synthesize participants’ interview responses supported by user log data to substantiate additional insights and suggest potential opportunities for future research directions. Our findings cover three key aspects: i) the utility of AI-personalized translation for scientific text, ii) the impact of interactive reading features on user experiences, and iii) the potential use cases and concerns.
5.1. Quantitative Findings from User Logs and Pre-study Surveys
5.1.1. Users’ Exploration Behaviors
Participants generated a total of 268 AI-personalized translations using TranSlider, with an average of 8.93 translations explored per session (min: 5, max: 19). We observed a positive correlation (Pearson’s $r=0.36$) between the degree of personalization and length of the translation (see Figure 3), showing that more personalized translations tended to be slightly longer. This subtle increase in length could be attributed to the need for integrating more detailed personal context in highly personalized translations, whereas low personalization required little to no incorporation of personal information.
Figure 3. The correlation between the personalization degree and personalized translation length (#words).
Participants demonstrated various exploration trajectories, including: i) incremental—starting from low degrees like 2 and moving toward higher degrees, ii) decremental—starting from high degrees like 98 or 100 and moving toward lower degrees, and iii) edge degrees—exploring the extremes (0 or 100) before exploring moderate degrees. Regardless of their exploration starting points, participants often revisited and compared previous translations, moving back and forth between different personalization degrees. While participants largely explored edge degrees more (0 and 100), many of them eventually favored translations with moderate degrees (see 4(a)). The mean value of participants’ favorite translations was 53.14 (SD: 33.24, Median: 51.0), indicating a tendency towards moderate personalization rather than extremes such as hyper-personalization or none at all.
5.1.2. Motivations for Reading Scientific Articles: Extrinsic and Intrinsic
We categorize participants’ motivation for scientific articles as extrinsic and intrinsic based on their pre-study interview responses. Those with extrinsic motivations (N=7) read scientific articles primarily when required for work-related purposes, while intrinsically motivated individuals (N=8) read them voluntarily, even in their spare time and regardless of work relevance. 4(b) illustrates the distribution of participants’ preferred translation degrees within each motivation category. Descriptively, we see that participants with extrinsic motivation tend to prefer a lower degree of personalization (M: 40.52, SD: 29.84), while those with intrinsic motivation favor higher degrees (M: 62.75, SD: 29.35). While our study was not designed to compare statistical differences between these two groups (Montgomery et al., 2018), this descriptive result shows that motivation could be an important factor in designing personalization interfaces.
(a) Participants tend to explore extreme personalization degrees (0 or 100) more frequently, yet their preferred translations typically fell within moderate ranges, despite less exploration of middle ones.
(b) Those with extrinsic motivation for reading scientific articles tend to prefer personalization lower degree, while those with intrinsic motivation liked translations with higher degrees.
Figure 4. (a) shows the histogram of two distributions: all translations vs. user-selected favorite translations. (b) displays a box plot of favorite translations by different motivations for reading scientific articles.
5.2. Utility of AI-Personalized Translation for Scientific Text (RQ1)
All participants reported the benefits of personalized translations for scientific articles. However, individuals preferred varying degrees of personalization for different reasons. We asked participants to explain their preferences and rationale with concrete examples of translations they had explored. Based on these examples, we categorized their responses into two groups: those preferring higher degrees of personalization (ranging from 51 to 100) and those preferring lower degrees (ranging from 0 to 45).
5.2.1. Participants Who Preferred Higher Degree Liked Relatable and Informative Translations
Participants who preferred translations with higher degrees of personalization appreciated that these highly relatable explanations facilitated their understanding of complex ideas (e.g., P7, P8, P9, P10, P14). They further stated that these personalized translations, which extensively incorporated participants’ personal contexts, resonated with them and were funny. For instance, P10 appreciated the translation began with ‘running’, one of her hobbies, making challenging scientific concepts more relatable and easier to understand. Similarly, P14 found some personalizations funny as they incorporated the safety and security aspects of his hometown and his cultural background.
P10: I like that it started with ‘running’, it brings it up here [in the translation]. But then it continued on the science [content]. Yeah, the science to me is usually the harder part. So I like that. It explained it [science] all at once, and then it started trying to simplify and use analogies.
Figure 5. Several examples of preferred translations. While each personalized translation is longer (two paragraphs), we present specific parts that participants pointed out to highlight why they liked or disliked the passage.
Furthermore, these participants felt that highly personalized translations offered detailed scientific content, including additional findings (P8) and detailed explanations of implications (P14).
P14: I feel like this last point was pretty good. Because it says, I’m not sure if the other ones really talked about it that much, but that the cleaner energy could be causing pollution. So the implication of the study is kind of good [useful].
These participants expressed dissatisfaction with translations of lower degrees. They felt that less personalized translations reduced the depth of scientific information (e.g., fewer science concepts and facts addressed in the original article). They also noted that lower degrees of personalization often employ analogies drawn from contexts that were intended to be more universally familiar to general audiences. However, these analogies turned out to be unfamiliar to many participants, ironically making it more challenging to comprehend the scientific articles (e.g., P9).
5.2.2. Participants Who Preferred Lower Degree Liked Concise Translations and Subtle Personalization
Some participants (e.g., P1, P2, P4, P5) preferred translations with lower degrees of personalization primarily because they demanded less mental load to comprehend. They emphasized the cognitive load required to process long personalized translations, stating that visibly longer translations immediately diminished their interest and willingness to read the material (e.g., P2, P5, P13). For instance, translations with lower degrees of personalization tend to be more concise (see Figure 3), which likely contributed to their reduced mental effort.
P5: I feel like I have to put in more effort to understand the analogy in order to understand the actual content. … it takes a long time, I think for me the length matters. I want it to be personalized, but I don’t want it lengthy.
P4: What I liked about less personal ones was that they felt more direct, versus this [high degree personalization] was like reading a lot of text to get to the main point. Yeah, the personalization makes the actual message longer.
They also appreciated that the analogies were subtly integrated into lower degrees, making them easy to understand and requiring minimal cognitive effort.
P5: If I’m ever reading a scientific article, the goal is not to read a lot, but [to figure out] mostly in the shortest amount of time, what is the most information I can take away. So I think having more subtle analogies allowed me to get to the understanding better, but still have the key takeaways.
5.2.3. Experimenting with Profile Edits
Despite acknowledging the benefits of personalized translations, regardless of their preferred degrees, participants reported that the translations were sometimes unnatural and difficult to understand. Specifically, they found that certain aspects of their background (personalization profile) were not appropriate to explain the given scientific content. Profile editing enabled them to address this issue by experimenting with differentprofiles. Participants enjoyed experimenting with their profiles to explore various personalization options. They found it entertaining and fun as they could tell that profile modifications directly influenced the translations. For example, P2 pretended to be a five-year-old Chinese boy, and P3 added “going to a bar” as a hobby to receive bar-related analogies.
P3: Bar one. I picked the right analogy… And I didn’t mean to. I was just like having fun with it. But that’s a really good analogy, for PFAS I would say.
Participants noted that hobby-related information influenced the analogies used in translations, while age settings impacted the overall tone.
P6: Oh, I thought it was pretty good. I noticed, specifically, when I changed the profile I liked seeing that drastic differentiation. As you mentioned, age was potentially a factor. I’m glad to see that [age] was taken into consideration in the tone of the message.
5.3. Impact of Interactive Reading Features on User Experience (RQ2)
We analyze participants’ responses to two features of the tool: generating multiple translations per article and slider interaction to control the personalization degree.
5.3.1. Compounding Understanding with Multiple Translations
Most participants reported several benefits of reading multiple translations, describing it as assembling a complete picture from partial pieces. They also mentioned that reading multiple translations enabled a cumulative understanding of complex scientific concepts iteratively (e.g., P1, P5, P7, P8, P10) and even helped correct initial misunderstandings (e.g., P2). For example,
P10: As I read more [translations], I understand the topic more. It was helpful as I kept going. Because I feel like I was picking up some small details here and there, and it pieced it together at the end. It definitely wasn’t distracting, like completing a big picture like by collecting more positive pieces.
Furthermore, through an iterative process of generating and reading multiple translations, P2 discovered inconsistencies in their initial understanding. This process helped them identify and correct conceptual misunderstandings that they had not initially recognized, ultimately leading to a more accurate comprehension of the material.
P2: I thought the sugar was BRCA2. It’s actually not. [I noticed from the later translations that] the protein is the BRCA2 and the sugar bypasses the BRCA2.
However, despite such benefits, a few participants (e.g., P2, P3) expressed reluctance to explore multiple translations in realistic scenarios, especially for reading multiple versions of scientific articles that did not align with their interests.
P3: In real life, I wouldn’t want to read the same thing in a different analogy. If the 1st one didn’t work, I’d be like, ‘Okay, whatever’ you know, I wouldn’t rerun it. I don’t have enough time. [On the other hand] I’ve read different political articles this year, especially with an election year. If there’s something I don’t understand, I’ll go find information elsewhere. In the same way, I would potentially rerun this same interface to see a different analogy or different versions [if I’m interested in the topic].
This shows that articles that deeply interest participants are a strong motivator for them to read multiple translations and gain a better understanding. While TranSlider could offer valuable benefits for comprehension, user engagement may vary depending on individual interests and the perceived relevance of the content.
5.3.2. Balancing Control vs. Automation Afforded by the Slider
Participants generally recognized and appreciated the gradient differences in personalized translations across the 0-100 spectrum on the slider control. All participants found the slider interaction to be simple, straightforward, and enjoyable, requiring minimal effort to modify the translations of scientific articles.
P15: I think that it gave me a lot of control, and I like the control [the agency] that it gave me. I’m selecting what I want to see, when the model is doing [translation] for you. I think that the more you can interact with it [through slider] and kind of control, it makes you feel welcome.
While participants agreed that the slider was simple and engaging, their desired level of control differed. Some participants expressed interest in having more granular control over the translations (e.g., P3, P7, P8, P9, P10), suggesting that the slider alone might not be sufficient. For example, they wanted to type specific numbers directly to modify the degrees of personalization in addition to slider interaction (e.g., P9) or to be more interactive like ChatGPT (e.g.,  (Laban et al., 2024)).
P10: It would be nice if there was an area of more personalization where I tell them what I like about it, and then they do that [personalization] more or do it less. Like, ‘I would like you [the AI model] to specify less, or I would like you to specify more on either analogy or on the science of it.’
In particular, P10 proposed a feature to allow users to highlight specific parts of the personalized translations and modify them instead of changing the entire translation. They envisioned that this feature could enable users to request more detailed elaboration or varying degrees of personalization for each selected part, providing more granular control over how the tool interprets and modifies different text parts.
In contrast, others preferred minimal control or even automated translations without user intervention at all. Several participants (e.g., P2, P13) found the slider feature overwhelming as it provides too many options (i.e., 100 different versions in the range 0 to 100). For example, P13 suggested 10 or even fewer options for personalization rather than a scale of 0-100. Some proposed having just two versions of the article: original and translated, with an on/off toggle.
Furthermore, P2 suggested a completely automated personalization system. He denied the need for any interactive control options. Rather than investing time or effort in steering various personalization degrees, he envisioned a future where AI would automatically deliver optimally personalized translations based on user preferences and article topics. He believed that AI systems should be capable of determining the optimal degree of personalization for each user and context without manual input. Interestingly, he did not worry about his privacy for such a system as he perceived clear benefits.
P2: Maybe that [the optimal degree of personalization] should be already part of your profile – how much personalization do you want. Then it’s like, I just want the optimal translation. It’s easier for someone to go on a blog and just have it [the optimally personalized translation] straight up there.
5.4. Potential Use Cases and Concerns
Participants identified the potential use cases and some concerns about this type of tool. Below we describe two primary use cases: learning unfamiliar content and tailoring their own ideas to different audiences. We then present the participants’ two main concerns: reliability of malleable translations and reduced access to the original source.
5.4.1. Learning Unfamiliar Content for Various Motivations (Use Case I)
Participants envisioned using this type of tool to learn unfamiliar topics that they were interested in (e.g., technology (P1), political pledges (P3), economic content (P10), how-to documents (P15)). Particularly, P10 expected that this tool could expand the boundaries of domains that are possible to learn about, as any unfamiliar topic could be explained in a relatable context.
P10: I tend to dodge articles about Econ or business. That’s not my strong suit, so I feel like I would be much more willing to read those. Because usually they’re filled with so much jargon, so like it now opens up a field for me of something that I’m interested in learning.
Furthermore, P3 and P9 envisioned using this tool when they would need to learn something about an unfamiliar topic (e.g., airline compensation policy documents (P3)), which they are not intrinsically motivated to learn, but need to know about (extrinsic motivation).
P9: Maybe when I’m trying to find an article that I’m not actually interested in, but I need to get specific information about the topic that I’m not too familiar with where I need the translation. That’s where I would see myself using something like this.
Overall, participants found this tool useful to learn about various unfamiliar topics that they wished to explore. Some of them noted that it could help them expand their knowledge into new domains. Moreover, even for topics they found less interesting, the tool could help them quickly grasp essential concepts and key takeaways.
5.4.2. Tailoring Ideas to Communicate to Different Audiences (Use Case II)
One participant (P1) emphasized that this tool could be useful not only to understand unfamiliar topics but also to learn how to communicate complex ideas to diverse audiences. For example, the different degrees of personalization could provide tailored translations of how to explain technical products or business ideas to general audiences. She envisioned saving multiple translations to her personal library and using them during business meetings.
P1: If I need to explain this topic to somebody who’s in the industry is very different from how I will explain this to somebody who’s not technical. So it’s nice to know how I should explain this to somebody who has some knowledge and how to explain it to somebody who has no knowledge like the board of directors for their company.
5.4.3. Reliability of Malleable Translations from AI (Concern I)
While most participants focused more on the benefits, one participant (P8) raised concerns about the reliability of AI-generated personalized translations. He argued that personalized translations can vary with each interaction, raising questions about their consistency and trustworthiness.
P8: If I read that I need to use the knowledge, it’s a little more concerning, like I might not trust it as much because it keeps changing. If I had to reference it, or something, I probably wouldn’t trust this. It [AI] might misunderstand something.
He further stated that this concern would motivate him to check out the original paper for complete and accurate understanding. While he would not fully rely on the translations in his professional context, he acknowledged that the tool would be useful in deciding whether to read the original document.
5.4.4. Reduced Access to Original Source Material (Concern II)
One participant (P5) raised a contrasting concern: she was worried that AI-personalized translations might discourage people from accessing original documents. People might rely on these personalized translations rather than engaging with the original material because personalized translations are easier to digest. This could lead to a dependence on AI-processed content at the expense of direct engagement with source materials.
P5: The downside is that nobody would probably read the full paper outside of people within that industry or field. But I don’t think that’s necessarily bad, because the alternative is you have more people understanding key concepts.
P5 summarized the trade-off associated with this tool. On one hand, it might lead to a larger number of people gaining superficial knowledge across a wider range of topics, without developing a deep understanding or motivation to explore further. On the other hand, without such a tool, it would be less accessible and fewer people would be exposed to or aware of this knowledge in the first place. While acknowledging the concern about shallow understanding, P5 also emphasized increased accessibility to more scientific information.
6. Discussion
Participants preferred varying degrees of personalized translations. Some preferred concise and less contextual translations (lower degree), while others preferred relatable and contextual ones (higher degree). Interactive reading features enabled them to compound their understanding and even correct initial misunderstandings. Participants noted that these benefits would be particularly valuable when reading topics they were interested in. On the other hand, some participants seeking quick takeaways found little value in exploring multiple translations using the slider. For them, the effort required to explore various translations outweighed the perceived benefits. Given these findings, we discuss several implications of our study in supporting communication in collaborative contexts, maintaining user agency for AI alignment, and the reliability of AI outputs.
6.1. Enhancing Communication by Surfacing Diverse Perspectives
Our findings showed that AI-personalized translations can help users better comprehend scientific content. However, participants perceived broader benefits. They found that personalized translations helped them grasp complex ideas from varied perspectives. The ability to explore multiple viewpoints can prove useful in contexts such as education and cross-domain collaboration.
Education has grown increasingly complex, particularly with the widespread adoption of online learning (Li et al., 2021). While online platforms provide greater scalability, instructors face persistent challenges in delivering effective instruction to large, diverse student cohorts (Kulkarni et al., 2015). AI-powered personalization tools can help address these challenges by offering instructors tailored translation suggestions adapted to different students. With appropriate privacy safeguards, these systems could utilize student profiles to recommend more effective concept explanations, enabling more relevant, contextualized interactions between students and instructors. Future research should explore the impact of these personalized translations on learning outcomes.
In cross-disciplinary contexts, effective communication presents similar challenges (Agapie et al., 2022). While cross-domain collaboration is invaluable, several obstacles impede clear communication, such as divergent terminology (Hou and Wang, 2017), varying data interpretations (Kim et al., 2022), and different motivations (Mao et al., 2019). Research shows that new ideas are better understood when presented through familiar examples (Kim et al., 2022). AI-personalized translation tools can bridge these gaps by adapting explanations to collaborators’ backgrounds and expertise, making complex concepts more accessible. This capability is particularly valuable in practical scenarios where cross-disciplinary collaboration is essential for solving complex problems (e.g.,  (Saha et al., 2021; Yoo et al., 2024)). Users can input their collaborators’ domain contexts into the system to receive guidance on effectively communicating their ideas in a way that resonates with their collaborators’ expertise and perspectives.
6.2. Maintaining User Agency in the Pursuit of AI Alignment
Traditionally, the machine learning community has focused on aligning AI models with human intentions through techniques like fine-tuning or reinforcement learning (Ouyang et al., 2022). While these approaches have been effective, they diminish the plurality of cultures and values represented among AI users (Agarwal et al., 2024; Anderson et al., 2024). In this paper, we present an alternative approach that emphasizes user agency. Rather than optimizing the model directly, we designed an interactive tool that empowers users to actively steer and align outputs according to their individual preferences.
This design intervention respects the diversity of experiences and backgrounds of intended general audiences, essential for good science communication (Nisbet and Scheufele, 2009). Our interface provides three key features to support user agency: a) a slider that spans a spectrum of personalization degrees, b) a profile editing function, and c) a history box. These features allow users to influence model outcomes without technical expertise, save personalized preferences, and iteratively explore various outputs until they find desirable results. This approach builds on extensive work in the HCI community which has demonstrated the value of intention-centric design across various domains (e.g.,  (Horvitz, 1999; Amershi et al., 2019)).
While highly accurate AI models might better serve diverse user needs, achieving such accuracy often requires extensive user data collection, raising concerns of cost, privacy, and selection bias amongst annotators (Sambasivan et al., 2021). Our design approach provides an alternative: rather than relying on AI-powered systems to perfectly model the plurality of user preferences, we empower users to actively align AI outputs to their individual preferences. This approach not only preserves user privacy but also maintains human agency in human-AI interaction, addressing critical concerns about AI overreliance in our society (Buçinca et al., 2021a; Okolo et al., 2024).
6.3. The Reliability of AI-Personalized Translations
Participants expressed contrasting concerns about the reliability of AI-generated translations for science communication. While some participants were reluctant to rely on the translations, others were worried that people might depend too heavily on them. As AI technology becomes more sophisticated, the higher accuracy of translations with fewer instances of incorrect or hallucinated content might mitigate such reliability concerns (Wu et al., 2017). However, regardless of translation quality, concerns about AI agencies and the trustworthiness of AI may persist. The fact that AI-generated translations are not human-authored continues to affect the perceived trust (Ryan, 2020). This might be due to the unclear responsibility; it is vague who bears the responsibility for mistranslations and their potential consequences. For instance, a participant mentioned that they would not trust AI translations for work-related purposes without verifying the original source, as they would be personally accountable for any errors in the work.
To mitigate risks, we may consider encouraging readers to verify translations against original source materials. However, this might be bothersome. As a participant noted, many people would overly rely on the AI translations and blindly trust what is generated simply because it is convenient, and as we noticed some people do not want to invest mental effort (Buçinca et al., 2021b). This balance between utilizing AI translations and encouraging verification of original sources presents a unique challenge, particularly given the current state of scientific communication. While AI can facilitate making scientific content more accessible, the verification process might be still infeasible as many original scientific articles use overly complex language and jargon, creating barriers for general audiences. Alternatively, we can consider authors (i.e., scientists) to verify several versions of AI-translations before publishing to the general public (Kim et al., 2024). For example, scientists can go through 5-10 different translations with varying degrees of personalization and approve a few of them as verified translations. This would remove the onus of verification from the reader while allowing them to receive the benefits of multiple translations. Further explorations are needed to find scalable approaches such that the verification by authors does not become cumbersome.
6.4. Limitations and Future Studies
Our study had several limitations. First, it was conducted in a controlled setting. Specifically, a small sample size of only 15 participants were asked to explore two pre-selected science articles from health and environment domains using our tool. Future studies should consider a large-scale in-the-wild study where participants could choose from a wider range of scientific articles. Secondly, our study setup required participants to provide feedback based on brief interactions with the tool. The long-term effects of interactive reading experiences with AI-personalized translation tools may differ from our findings. Future longitudinal studies are warranted to provide a more comprehensive understanding of how users interact with and benefit from these types of tools over extended periods. Lastly, our participant demographics were limited, with most participants (12 out of 15) being in their twenties and recent college graduates. Only three participants were from other age groups. This age-homogeneous participant pool may have influenced the findings of our study. Future studies should include participants from diverse backgrounds in terms of age, education level, and culture to capture a broader range of perspectives on AI-personalized translation tools. For instance, the impact of AI-personalized translation tools in high school science learning classes (Lee and Perret, 2022) or children-parent learning contexts (Cingel and Piper, 2017) could be different from our findings.
7. Conclusion
Our study investigated the potential of AI-personalized translations to enhance science communication in digital media platforms. We developed TranSlider, an AI-powered reading interface that generates personalized translations of scientific text for general audiences. The tool’s key feature involves an interactive slider that allows users to adjust the degree of personalization to generate tailored translations. Through an exploratory study with 15 participants, we found that participants preferred varying degrees of personalization for the tailored translations, from deeply relatable and contextualized translations to short and slightly contextualized. We also found that reading multiple translations enabled them to have a compounding understanding of scientific content. We present several implications in supporting communication across diverse backgrounds in collaborative contexts and highlight the importance of user agency in AI-powered systems to facilitate desirable outputs.